OpenStudy (anonymous):

I need 2x * x to = 24. Anyone know what x is?

OpenStudy (anonymous):

\[+2\sqrt{3}\]

OpenStudy (anonymous):

ignore the previous one\[\pm 2\sqrt{3}\]

OpenStudy (anonymous):

whats that symbol mean?

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

the + with a line under it

OpenStudy (anonymous):

that is a sign of plus and minus together

OpenStudy (anonymous):

so what do i do with it?

OpenStudy (anonymous):

it means x can take 2 values that are \[+2\sqrt{3}\] and \[-2\sqrt{3}\]

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

welcome!! :)